# Unveiling Linguistic Regions in Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated considerable cross-lingual alignment and generalization ability. Current research primarily focuses on improving LLMs’ cross-lingual generalization capabilities. However, there is still a lack of research on the intrinsic mechanisms of how LLMs achieve cross-lingual alignment. From the perspective of region partitioning, this paper conducts several investigations on the linguistic competence of LLMs. We discover a core region in LLMs that corresponds to linguistic competence, accounting for approximately 1% of the total model parameters. Removing this core region by setting parameters to zero results in a significant performance decrease across 30 different languages. Furthermore, this core region exhibits significant dimensional dependency, perturbations to even a single parameter on specific dimensions leading to a loss of linguistic competence. Moreover, we discover that distinct monolingual regions exist for different languages, and disruption to these specific regions substantially reduces the LLMs’ proficiency in those corresponding languages. Our research also indicates that freezing the core linguistic region during further pre-training can mitigate the issue of catastrophic forgetting (CF), a common phenomenon observed during further pre-training of LLMs. Overall, exploring the LLMs’ functional regions provides insights into the foundation of their intelligence.
@@ 1 Introduction
Q1: Does a core linguistic region exist within LLMs that facilitates cross-lingual generalization? By conducting further pre-training across six languages and evaluating models’ parameter importance (Section 2.2), we discover a core region in LLMs corresponding to linguistic competence, which accounts for approximately 1% of the model’s total parameters. As shown at the top of Figure 1, removing this region (setting parameters to zero) consistently leads to a significant decline in performance across 30 test languages (Section 3.2).
Furthermore, by visualizing the core linguistic region (Figure 2), we observe that the linguistic core region of LLMs exhibits significant dimensional dependency. In certain dimensions, only perturbing a single parameter could lead to the model losing its linguistic competence (Section 3.3). Additionally, ablation study in 3.3 shows that beyond outlier dimensions, other non-outlier dimensions in this region are also critical.
@@ 3.2 Core Linguistic Competence Region
In this section, we conduct further pre-training experiments on LLaMA-2 across six languages, aiming to explore and identify the core linguistic region in LLMs. Here, we define it as "core linguistic", attributed to its minimal proportion within LLMs, constituting only 1% of the total parameters, and its association with model’s linguistic competence modeling.
@@
In the experiment corresponding to Table 2, we initially zero out various regions within LLaMA. Consistent with the findings from Table 1, removing the ‘Top’ region leads to a loss of linguistic competence, whereas the ‘Bottom’ region don’t. However, in this experiment, we sought to ascertain if LLaMA could reacquire its lost cross-lingual generalization competence. Thus, we train on different amounts of Chinese Zhihu corpus and evaluate on Chinese Wechat and English Falcon corpora. The results indicate that unlike the ‘Bottom’ region, if the ‘Top’ region is removed and frozen, the model have to relearn basic language rules in other regions based on the provided Chinese Zhihu corpus, but these rules are inherently biased towards Chinese. Consequently, while its proficiency in Chinese is restored, the English perplexity remains high (). If the ‘Top’ region is removed but not frozen, the model can rebuild its linguistic competence in-place. As its proficiency in Chinese is restored, so is its proficiency in English. This implies that the ‘Top’ region encodes generalizable fundamental linguistic competence. When ‘Top’ region is zeroed-out and frozen, other regions difficultly adapt to regain core linguistic competence.
@@ 3.3 Dimensional Dependence
Instead of discretely removing different unstructured parameters, we selectively remove structured certain rows or columns for each matrix, especially those dimensions encompassing a significant number of ‘Top’ region parameters, termed as ‘Top’ dimensions. As illustrated in Table 3, we attempt to remove the columns of FFN.down and Attn.k/q/v, as well as the rows of Attn.o. The results indicate that removing just these ‘Top’ dimensions leads to a substantial decline in the model’s linguistic competence. However, disturbances to the ‘Middle’ ,‘Bottom’ and ‘Random’ dimensions do not yield noticeable effects. Selecting the dimensional region only from the Attention matrix or inverting rows and columns removals lead to similar findings, as described in C.
@@
Output Under Perturbations
To visually illustrate the impact of the linguistic competence region on the model’s output quality, we use "Fudan University is located in" as a premise and observe the model’s outputs under different parameter perturbations. The results are shown in Figure 4. Compared to randomly selected 4200-th dimension, perturbing model on 2100-th dimension significantly leads to model loses its linguistic competence, producing error or even nonsensical strings.
@@
Considering that the collapse of PPL may be possibly caused solely by the removal of outlier dimension parameters, rather than the collective effect of the entire linguistic region, we conduct an ablation experiment without removing the outlier dimensions’ parameters (1512/2533 for LLaMA-2-7B and 2100/4743 for LLaMA-2-13B). The results for the 13B model, presented in Table 4, indicate that while outlier dimensions are contained within the core linguistic region, non-outlier dimensions in this region are also critical. Not altering any row or column of the matrix parameters for the two outlier dimensions also results in a great increase in PPL, although the collapse of PPL is slowed. This finding reveals that all dimensions in the core linguistic region are tightly interrelated, and disrupting even a small part of it can lead to a PPL collapse. For a more detailed analysis of the experiments and results for the 7B model, please refer to Appendix E.
@@
| Attn.o(row) Attn.k/q/v(column) FFN.down(coulumn) | 1 | 5218.1 | 11.079 |
@@ 3.4 Monolingual Region
In this section, we wonder if LLMs possess distinct regions within an individual language. Unlike the core linguistic regions described in Section 3.2, a monolingual family region only has a strong correlation with certain languages, and removing it will only cause significant influence on LLMs’ proficiency in those corresponding languages.
@@
Different from Section 3.2, we initially identify and select the 1% ‘Top’ and ‘Bottom’ regions for each of the six languages (Arabic, Spanish, Russian, Chinese, Korean, and Vietnamese) according to Equation 4, then deduplicate these regions. For the target language region, we exclude any regions that overlap with the ‘Top’ and ‘Bottom’ regions of the other five languages, aiming to eliminate the core regions and critical dimension corresponding to the model’s fundamental linguistic abilities. We denote , and as the total set of six languages and the ‘Top/Bottom’ regions before and after deduplication, respectively. Language ’s own region is computed as follows:
In Appendix F, we visualize the distribution Attn.q matrix in ‘Arabic’ and ‘Vietnamese’ regions and discover minimal overlap between them.
@@
Unlike removing core regions or dimensions in Section 3.2 and 3.3, we discover that removing monolingual regions will only significantly affect the ability of the target languages and their closely related languages with similar letter elements or sentence structure. For example, if we remove only the region for Russian alone, selected from (10K) or (100K) samples respectively, as shown in Table 5, only Russian itself and Ukrainian have significant increases in PPL when removing ‘Top’ region. We speculate this to the fact that Russian and Ukrainian are relatively similar in terms of sentence structure and constituents, both belonging to the Slavic group. A similar phenomenon is observed if removals are changed to the regions for each of other five languages, see Appendix F for more details.
@@
1) The average logits probability for option ‘A’ on the ArabicMMLU evaluation increases from 0.36 to 0.64. 2) For the evaluation where the correct answer is one of the options ‘B/C/D/E’ in ArabicMMLU, the model’s accuracy decreases dramatically from 25.6% to just 1.5% when the ‘Arabic’ region is removed. However, if the ‘Vietnamese’ region is removed, the model’s accuracy remains almost unaffected at 26.7%. 3) For the MMLU evaluation, the model’s accuracy is minimally impacted by the removal of monolingual regions. Compared to a baseline accuracy of 42.46%, the removals of the ‘Arabic’ and ‘Vietnamese’ regions result in accuracies of 39.27% and 39.68%, respectively.
Furthermore, using "There are 365 days in a year and 12" as a prompt for generation, we test outputs on the removal of the removing monolingual regions. More details are in Appendix F.
@@ 3.5 Further Pre-training Optimization
In this section, we demonstrate that stabilizing the core linguistic regions identified in Section 3.2 mitigates the catastrophic forgetting (CF) issue (McCloskey and Cohen, 1989; Kemker et al., 2018) in LLMs, while maintaining learning proficiency comparable to full-scale fine-tuning in target language. Our experimental setup involves further pre-training LLaMA-2-7B on 100,000 Arabic sentences, with a batch size of , a maximum token count of , and learning rates () and or , employing perplexity (PPL) as the evaluation criterion.
@@
Traditional full-scale fine-tuning, when increasing the learning rate or the amount of corpus data, enhances learning in the target language but aggravates forgetting in non-target languages. As shown in Table 2, since LLaMA-2 is primarily trained on English corpora, conducting a second stage of pretraining solely on large-scale Chinese corpora can lead to the forgetting of English competence. Additionally, as depicted on the left side of Figure 6 in blue line, increasing from (dotted line) to (solid line) under full-scale fine-tuning boosts the acquisition of the target language (Arabic), while simultaneously accelerates the forgetting rate of the non-target languages (English and Chinese), as shown in the middle and right side.
@@
We hypothesize that CF problem occurs due to the amplification of parameter adjustments when increasing the learning rate, which leads to significant shifts in the core linguistic region, adversely affecting language alignment. To mitigate this, we protect the core linguistic region and key dimensions by freezing the ‘Top 5%’ core language area for fine-tuning, as shown by the red line in Figure 6. At a of (dotted line), the difference between freezing fine-tuning and full-scale fine-tuning is minimal. However, when the increases to (solid line), freezing fine-tuning not only facilitates faster learning in the target language (preserving comparable performance in Arabic PPL: 3.557 vs. 3.566), but also significantly reduces the forgetting of non-target languages (showing improvements in English and Chinese PPL: 18.796 vs. 20.557 and 90.84 vs. 563.423, respectively).
The potential reason for this phenomenon may lie in the preservation of the core regions within the cross-lingual alignment competence. Restricting the magnitude of updates in the core region’s parameters is a future strategy we intend to employ. Notably, unlike regularization methods (Srivastava et al., 2014; Goodfellow et al., 2014), such approaches restricts to a minimal core region in LLMs, and can be implemented alongside blending previous data, retraining the entire network, or possibly only the final layers, without adding additional components to the model.
4 Realated Work
@@
Prior works aimed to extract a sub-network capable of executing specific downstream tasks (Frankle and Carbin, 2019; Zhang et al., 2022) or task-specific subspaces to limit parameter fine-tuning within it (Aghajanyan et al., 2021; Zhang et al., 2023). For parameter importance estimation, an effective metric is to use parameter magnitude (Zhu and Gupta, 2018; Renda et al., 2020; Zafrir et al., 2021). Another metric involves estimating the sensitivity of parameters (Molchanov et al., 2019; Sanh et al., 2020; Liang et al., 2021; Sapkota and Bhattarai, 2023). In this work, we employ the latter method to select the most crucial parameters to unveil linguistic regions. (Frankle and Carbin, 2019) extracted a complete sub-network for downstream tasks, as exemplified by the optimal scale of the lottery network at 21.2%. However, our findings indicate that the linguistic regions is a much smaller (1%) functional region.
@@
Prior works have shown that multilingual LMs rely on a shared subword vocabulary and joint pre-training across multiple languages (Pires et al., 2019; Cahyawijaya et al., 2023; Wu and Dredze, 2019). However, new insights highlight these models’ capacity for learning universal semantic abstractions (Artetxe et al., 2020; Chi et al., 2020) and demonstrate that mBERT (Devlin et al., 2019) embeddings of similar words in similar sentences across languages are approximately aligned already (Cao et al., 2020; Conneau et al., 2020; Xu et al., 2022). Analysis from a hierarchical perspective reveals that classifiers linked to different BERT (Devlin et al., 2019) layers assess semantic features through varied probe tasks (Lin et al., 2019; Jawahar et al., 2019). In this work, we introduce a parameter partitioning perspective within LLMs, identifying core linguistic and monolingual regions, which underpin cross-lingual alignment and language-specific characteristics, respectively.
@@
Multiple studies have identified outlier dimensions in Transformer-based LMs (Kovaleva et al., 2021; Dettmers et al., 2022). Researches have found that certain outlier dimensions in PLMs are highly sensitive to the fine-tuning of downstream tasks (Kovaleva et al., 2021; Puccetti et al., 2022). Furthermore, (Rudman et al., 2023; Puccetti et al., 2022) discovered that these outlier dimensions encode task-specific knowledge, and disabling these dimensions significantly degrades model performance. (Luo et al., 2021) attribute positional embeddings to the emergence of outliers, while (Rajaee and Pilehvar, 2022) reported inconsistent results. Additionally, (Sun et al., 2024) found that these outlier dimensions exhibit significantly larger activation values than others in LLMs. Our findings in Section 3.3 demonstrate that beyond the outlier dimensions, other non-outlier dimensions within the core linguistic region also play an indispensable role in the model’s core linguistic competence.

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated considerable cross-lingual alignment and generalization ability. Current research primarily focuses on improving LLMs’ cross-lingual generalization capabilities. However, there is still a lack of research on the intrinsic mechanisms of how LLMs achieve cross-lingual alignment. From the perspective of region partitioning, this paper conducts several investigations on the linguistic competence of LLMs. We discover a core region in LLMs that corresponds to linguistic competence, accounting for approximately 1% of the total model parameters. Removing this core region by setting parameters to zero results in a significant performance decrease across 30 different languages. Furthermore, this core region exhibits significant dimensional dependence, perturbations to even a single parameter on specific dimensions leading to a loss of linguistic competence. Moreover, we discover that distinct monolingual regions exist for different languages, and disruption to these specific regions substantially reduces the LLMs’ proficiency in those corresponding languages. Our research also indicates that freezing the core linguistic region during further pre-training can mitigate the issue of catastrophic forgetting (CF), a common phenomenon observed during further pre-training of LLMs. Overall, exploring the LLMs’ functional regions provides insights into the foundation of their intelligence 111 Our code is released in https://github.com/zzhang0179/Unveiling-Linguistic-Regions-in-LLMs..
@@ 1 Introduction
Q1: Does a core linguistic region exist within LLMs that facilitates cross-lingual alignment and generalization? By conducting further pre-training across six languages and evaluating models’ parameter importance (Section 2.2), we discover a region in LLMs corresponding to the core linguistic competence, which accounts for approximately 1% of the model’s total parameters. As shown at the top of Figure 1, removing this region (setting parameters to zero) consistently leads to a significant decline in performance across 30 test languages (Section 3.2).
Furthermore, by visualizing the core linguistic region (Figure 2), we observe that the linguistic core region of LLMs exhibits significant dimensional dependence. In certain dimensions, only perturbing a single parameter could lead to the model losing its linguistic competence (Section 3.3). Additionally, ablation study in 3.3 shows that beyond outlier dimensions, other non-outlier dimensions in this region are also critical.
@@ 3.2 Core Linguistic Competence Region
In this section, we conduct further pre-training experiments on LLaMA-2 across six languages, aiming to explore and identify the core linguistic region in LLMs. Here, we define the region as "core linguistic", attributed to its minimal proportion within LLMs, constituting only 1% of the total parameters, and its association with model’s linguistic competence modeling.
@@
In the experiment corresponding to Table 2, we initially zero out various regions within LLaMA. Consistent with the findings from Table 1, removing the ‘Top’ region leads to a loss of linguistic competence, whereas the ‘Bottom’ region don’t. However, in this experiment, we sought to ascertain if LLaMA could reacquire its lost cross-lingual generalization competence. Thus, we train on different amounts of Chinese Zhihu corpus and evaluate on Chinese Wechat and English Falcon corpora. The results indicate that unlike the ‘Bottom’ region, if the ‘Top’ region is removed and frozen, the model have to relearn basic language rules in other regions based on the provided Chinese Zhihu corpus, but these rules are inherently biased towards Chinese. Consequently, while its proficiency in Chinese is restored, the English perplexity remains high (). If the ‘Top’ region is removed but not frozen, the model can rebuild its linguistic competence in-place. As its proficiency in Chinese is restored, so is its proficiency in English.
This implies that the ‘Top’ region encodes multilingual linguistic competence. When ‘Top’ region is zeroed-out and frozen, other regions difficultly adapt to regain the core linguistic competence.
@@ 3.3 Dimensional Dependence
Instead of discretely removing different unstructured parameters, we selectively remove structured certain rows or columns for each matrix, especially those dimensions encompassing a significant number of ‘Top’ region parameters, termed as ‘Top’ dimensions. As illustrated in Table 3, we attempt to remove the columns of FFN.down and Attn.k/q/v, as well as the rows of Attn.o. The results indicate that removing just these ‘Top’ dimensions leads to a substantial decline in the model’s linguistic competence. However, removals to the ‘Middle’ ,‘Bottom’ and ‘Random’ dimensions do not yield noticeable effects. Selecting the dimensional region only from the Attention matrix or inverting rows and columns removals lead to similar findings, as described in C.
@@
Considering that the collapse of PPL may be possibly caused solely by the removal of outlier dimensions, rather than the collective effect of the entire linguistic region, we conduct an ablation experiment without removing the parameters of outlier dimensions (1512/2533 for LLaMA-2-7B and 2100/4743 for LLaMA-2-13B). The results for the 13B model, presented in Table 4, indicate that while outlier dimensions are contained within the core linguistic region, non-outlier dimensions in this region are also critical. Not altering any row or column of the matrix parameters for the two outlier dimensions also results in a great increase in PPL, although the collapse of PPL is slowed. This finding reveals that all dimensions in the core linguistic region are tightly interrelated, and disrupting even a small part of it can lead to a PPL collapse. For a more detailed analysis of the experiments and results for the 7B model, please refer to Appendix E.
Output Under Perturbation
To illustrate the impact of the linguistic competence region on the model’s output quality, we use "Fudan University is located in" as a prompt input and observe the model’s outputs under different parameter perturbations. The results are shown in Figure 4. Compared to randomly selected 4200-th dimension, perturbing model on 2100-th dimension significantly leads to model loses its linguistic competence, producing error or even nonsensical strings.
@@
| Attn.o(row) Attn.k/q/v(column) FFN.down(column) | 1 | 5218.1 | 11.079 |
@@ 3.4 Monolingual Region
In this section, we wonder if LLMs possess distinct regions within different individual languages. Unlike the core linguistic regions described in Section 3.2, a monolingual region only has a strong correlation with certain languages, and removing it will only cause significant influence on LLMs’ proficiency in those corresponding languages.
@@
Different from Section 3.2, we initially identify and select the 1% ‘Top’ and ‘Bottom’ regions for each of the six languages (Arabic, Spanish, Russian, Chinese, Korean, and Vietnamese) according to Equation 4, then deduplicate these regions. For the target language region, we exclude any regions that overlap with the ‘Top’ and ‘Bottom’ regions of the other five languages, aiming to eliminate the core regions and critical dimension corresponding to the model’s fundamental linguistic competence. We denote , and as the total set of six languages and the ‘Top/Bottom’ regions before and after deduplication, respectively. Language ’s own region is computed as follows:
In Appendix F, we visualize the distribution of Attn.q matrix in ‘Arabic’ and ‘Vietnamese’ regions and discover minimal overlap between them.
@@
Unlike removing core regions or dimensions in Section 3.2 and 3.3, we discover that removing monolingual regions will only significantly affect the ability of the target languages and their closely related languages with similar letter elements or sentence structure. For example, if we remove only the region for Russian alone, selected from or samples respectively, as shown in Table 5, only Russian itself and Ukrainian have significant increases in PPL when removing ‘Top’ region. We speculate this to the fact that Russian and Ukrainian are relatively similar in terms of sentence structure and constituents, both belonging to the Slavic group. A similar phenomenon is observed if removals are changed to the regions for each of other five languages, see Appendix F for more details.
@@
1) The average probability for option ‘A’ on the ArabicMMLU evaluation increases from 0.36 to 0.64 if remove the ‘Arabic’ region. 2) As illustrated in Figure 5, the model’s accuracy in the ArabicMMLU (filtered) evaluation, where the correct answer is one of the options ‘B/C/D/E’, drops significantly from 25.6% to merely 1.5% when the ‘Arabic’ region is excluded. Conversely, the removal of the ‘Vietnamese’ region has a negligible impact on accuracy, which remains at 26.7%. 3) For the MMLU evaluation, the model’s accuracy is minimally impacted by the removal of monolingual regions. Compared to a baseline accuracy of 42.46%, the removals of the ‘Arabic’ and ‘Vietnamese’ regions result in accuracies of 39.27% and 39.68%, respectively.
Furthermore, using "There are 365 days in a year and 12" as a prompt for generation, we test outputs on the removal of the ‘Arabic’ and ‘Vietnamese’ regions. More details are provided in Appendix F.
@@ 3.5 Further Pre-training Optimization
In this section, we demonstrate that stabilizing the core linguistic regions (identified in Section 3.2) during further pre-training mitigates the catastrophic forgetting (CF) issue (McCloskey and Cohen, 1989; Kemker et al., 2018) in LLMs, while maintaining learning proficiency comparable to full-scale fine-tuning in target language. Our experimental setup involves further pre-training LLaMA-2-7B on 100,000 Arabic sentences, with a batch size of , a maximum token length of , and learning rates () of or , employing perplexity (PPL) as evaluation criterion.
@@
Traditional full-scale fine-tuning, when increasing the learning rate or the amount of corpus data, enhances learning in the target language but aggravates forgetting in non-target languages. As shown in Table 2, since LLaMA-2 is primarily trained on English corpora, conducting a second stage of pre-training solely on large-scale Chinese corpora can lead to the forgetting of English competence. Additionally, as depicted on the left side of Figure 6 in blue line, increasing from (dotted line) to (solid line) under full-scale fine-tuning boosts the acquisition of the target language (Arabic), while simultaneously accelerates the forgetting rate of the non-target languages (English and Chinese), as shown in the middle and right side.
@@
We hypothesize that CF problem occurs due to the amplification of parameter adjustments when increasing the learning rate, which leads to significant shifts in the core linguistic region, adversely affecting language alignment. To mitigate this, we protect the core linguistic region and key dimensions by freezing the ‘Top 5%’ core language area for fine-tuning, as shown by the red line in Figure 6. At a of (dotted line), the difference between freezing fine-tuning and full-scale fine-tuning is minimal. However, when the increases to (solid line), freezing fine-tuning not only similarly facilitates faster learning in the target language (preserving comparable performance in Arabic PPL: 3.557 vs. 3.566), but also significantly reduces the forgetting of non-target languages (showing improvements in English and Chinese PPL: 18.796 vs. 20.557 and 90.84 vs. 563.423, respectively).
The potential reason for this phenomenon may lie in the preservation of the core regions within the cross-lingual alignment competence. Restricting the magnitude of updates in the core region’s parameters is a future strategy that we intend to employ. Notably, unlike regularization methods (Srivastava et al., 2014; Goodfellow et al., 2014), such approaches restricts to a minimal core region in LLMs, and can be implemented alongside blending previous data, retraining the entire network, or possibly only the final layers, without adding additional components to the model.
4 Related Work
@@
Prior works aimed to extract a sub-network capable of executing specific downstream tasks (Frankle and Carbin, 2019; Zhang et al., 2022) or task-specific subspaces to limit parameter fine-tuning within it (Aghajanyan et al., 2021; Zhang et al., 2023). For parameter importance estimation, an effective metric is to use parameter magnitude (Zhu and Gupta, 2018; Renda et al., 2020; Zafrir et al., 2021). Another metric involves estimating the sensitivity of parameters (Molchanov et al., 2019; Sanh et al., 2020; Liang et al., 2021; Sapkota and Bhattarai, 2023). In this work, we employ the latter method to select the most crucial parameters to unveil linguistic regions. Unlike (Frankle and Carbin, 2019), which extracted a complete sub-network for downstream tasks, as exemplified by the optimal scale of the lottery network at 21.2%, our findings indicate that the linguistic region is a much smaller (1%) functional region.
@@
Prior works have shown that multilingual LMs rely on a shared subword vocabulary and joint pre-training across multiple languages (Wu and Dredze, 2019; Pires et al., 2019; Cahyawijaya et al., 2023). However, new insights highlight these models’ capacity for learning universal semantic abstractions (Artetxe et al., 2020; Chi et al., 2020) and demonstrate that embeddings of similar words in similar sentences across languages are approximately aligned already (Cao et al., 2020; Conneau et al., 2020; Xu et al., 2022). Analysis from a hierarchical perspective reveals that classifiers linked to different BERT (Devlin et al., 2019) layers assess semantic features through varied probe tasks (Lin et al., 2019; Jawahar et al., 2019). In this work, we introduce a parameter partitioning perspective within LLMs, identifying core linguistic and monolingual regions, which underpin cross-lingual alignment and language-specific characteristics, respectively.
@@
Multiple studies have identified outlier dimensions in Transformer-based LMs (Kovaleva et al., 2021; Dettmers et al., 2022). Researches have found that certain outlier dimensions in pre-trained LMs are highly sensitive to the fine-tuning of downstream tasks (Kovaleva et al., 2021; Puccetti et al., 2022). Furthermore, (Puccetti et al., 2022; Rudman et al., 2023) discovered that these outlier dimensions encode task-specific knowledge, and disabling these dimensions significantly degrades model performance. (Luo et al., 2021) attributed positional embeddings to the emergence of outliers, while (Rajaee and Pilehvar, 2022) reported inconsistent results. Additionally, (Sun et al., 2024) found that these outlier dimensions exhibit significantly larger activation values than others in LLMs. Our findings demonstrate that beyond the outlier dimensions, other non-outlier dimensions within the core linguistic region also play an indispensable role in the model’s core linguistic competence.